OpenStudy (anonymous):

Euler 's Cauchy equation Find general solution for 1) 4x2y” + 4xy’-y= 0 2) 2x2y” + 4xy’-5y= 0 Solve IVP for 1) x2y” - 4xy’+6y= 0 , y(1) = 1 ,y’(1) =0 2) x2y” +3xy’+ y= 0 , y(1) = 4 ,y’(1) =-2

OpenStudy (unklerhaukus):

\[x^2\frac{\mathrm d^2y}{\mathrm dx^2}+\alpha x\frac{\mathrm dy}{\mathrm dx}+by=f(x)\] substituting \(x=e^z\) \[\qquad\Downarrow\qquad\Downarrow\qquad\Downarrow\] \[\frac{\mathrm d^2y}{\mathrm dz^2}+(\alpha-1)\frac{\mathrm dy}{\mathrm dz}+by=f(e^z)\]

OpenStudy (anonymous):

to Find general solution for 1) 4x2y” + 4xy’-y= 0 as in this differential equcation degree of coefficient=order of derivative and it fit to form ax^2y''+bxy'+cy=0 ( so we use cauchy euler equcation) by comparing both eq we have a=4 ,b=4 ,c=-1 in cauchy euler equcation we use equcation am(m-1)+bm+c=0 putting values of a,b and c 4m(m-1)+4m-1=0 4m^2-4m-1=0 4m^2=1 m=1/2 and m=-1/2 as solution for m are real ans distint so general solution is in form y=c1x^(1/2)+c2x(-1/2)

OpenStudy (anonymous):

Solve IVP for 1) x2y” - 4xy’+6y= 0 , y(1) = 1 ,y’(1) =0 a=1 ,b=-4 ,c=6 so m(m-1)-4m+c=0 m^2-5m+6=0 (m-2)(m-3)=0 m=2,m=3 two different real root of m so general solution y=c1x^2+c2x^3 now we find particular solution by using iv as y(1)=1 so y=c1+c2 (1 eq) as y'(1)=0 by taking derivative of y=c1x^2+c2x^3 y'=2c1x+3c2x^2 0=2c1+2c2 (2 eq) by solving eq 1 and 2 we got c1=3 c2=-2 so particular solution y=3x^2-2x^3